OpenStudy (anonymous):

1. A company produces two varieties of pens: alpha and beta. Each alpha pen needs twice as much labour time as the beta pen. If only beta pens are manufactured, then the company can produce 500 pens per day. The market can take only up to 150 alpha pens and 250 beta pens per day from the company. The profits earned by the company on selling an alpha pen and a beta pen are Rs 8 and Rs 5 respectively. What are the respective numbers of alpha and beta pens that the company needs to produce in order to maximize the profit? A. (125,250) B. (150,200) C. (250,0) D. (250,250)

OpenStudy (anonymous):

help me guys

OpenStudy (anonymous):

Each alpha pen needs twice as much labour time as the beta pen. If only beta pens are manufactured, then the company can produce 500 pens per day. Therefore if only alpha pens were manufactured, then the company can produce 250 pens per day. If half the day was spent making ALPHA pens, how many could be made? If half the day was spent making BETA pens, how many could be made?

OpenStudy (ganpat):

x labors cost for beta 2x labor cost for alpha.. Market demand is 150 alpha + 250 beta.. profit for alpa = 8, while for beta = 5, since thats direct profit word, it would be including labor and other Expenses For Maximum profit ; take the solution and check out.. it would be from 1st two options.. 1. 125 & 250 ; Profit = 8 (125) + 250 (5) = 2250 2. 150 & 200 ; Profit = 8 (150) + 200 (5) = 2200 So option A...